OpenStudy (anonymous):

The acceleration experienced by a boat after the engine is cut off, is given by (dv/dt)=-kv^3,where k is a constant. if ''v(1)'' is magnitude of the velocity at cut off, find the magnitude of the velocity at time 't' after the cut off.

OpenStudy (kropot72):

\[\frac{dv}{dt}=-kv ^{3}=f(v)\] \[\frac{dv}{f(v)}=dt\] \[\int\limits_{}^{}\frac{dv}{f(v)}=t+C\] \[\int\limits_{}^{}\frac{1}{-kv ^{3}}dv=t+C\] \[-\frac{1}{k} \times -\frac{v ^{-2}}{2}=t+C\] \[\frac{1}{2kv ^{2}}=t+C\ ..................(1)\] \[v=v _{1}\ when\ t=0\] Substituting into (1) we get: \[\frac{1}{2kv _{1}^{2}}=C\ .................(2)\] Subtracting (2) from (1) we get: \[\frac{1}{2kv ^{2}}-\frac{1}{2kv _{1}^{2}}=t\ ................(3)\]

OpenStudy (kropot72):

Dividing both sides of (3) by t, we get: \[\frac{1}{2ktv ^{2}}-\frac{1}{2ktv _{1}^{2}}=1\] Rearranging gives: \[\frac{1}{2ktv ^{2}}=1+\frac{1}{2ktv _{1}^{2}}\] Multiplying both sides by 2kt we get: \[\frac{1}{v ^{2}}=2kt+\frac{1}{v _{1}^{2}}=\frac{1+2ktv _{1}^{2}}{v _{1}^{2}}\] Inverting both sides gives: \[v ^{2}=\frac{v _{1}^{2}}{1+2ktv _{1}^{2}}\] Taking the square root of both sides gives: \[v=\frac{v _{1}}{\sqrt{1+2ktv _{1}^{2}}}\]